### v4

# On Linear Quadratic Potential Games

Abstract
Our paper addresses characterizing conditions for linear quadratic (LQ) games to be potential. The desired properties of potential games in finite action settings, such as convergence of learning dynamics to Nash equilibria and challenges of learning Nash equilibria in continuous state and action settings, motivate us to characterize LQ potential games. Our first contribution is to consider two-player LQ games with full-state feedback and scalar states and action space, and to analytically verify that the set of potential games within this example is limited, essentially differing only slightly from an identical interest game. Given this finding, we restrict the class of LQ games to those with decoupled dynamics and decoupled state linear feedback information structure. For this subclass, we show that the set of potential games strictly includes non-identical interest games and characterize conditions for LQ games in this subclass to be potential. We further derive their corresponding potential function and prove the existence of a Nash equilibrium. Meanwhile, we highlight the challenges in characterizations of the Nash equilibrium for this class of potential LQ games, theoretically and through simulation.
keywords:
Distributed control; Dynamic games; Reinforcement learning in control.
thanks: This paper appears on Automatica
https://doi.org/10.1016/j.automatica.2025.112643
1 Introduction
Numerous emerging networked dynamic systems, such as smart grids and autonomous vehicles, rely on the decision-making of multiple interacting agents. Due to the increasing complexities of the dynamics and uncertainties of these systems, data-driven control has gained increasing attention. Reinforcement learning approaches like policy gradient methods have been widely applied in single-agent settings, and their convergence properties have been theoretically investigated [8]. However, multi-agent reinforcement learning has been less understood in both theory and practice.
Non-cooperative linear quadratic (LQ) games serve as a fundamental setting for understanding the convergence of multi-agent reinforcement learning algorithms. This class of games admits Nash equilibria in linear policies, which can be computed by solving a set of coupled Riccati equations [4]. Given these well-understood properties of LQ games, several recent studies have sought to theoretically understand multi-agent reinforcement learning algorithms within the LQ game setting.
Most prior works on learning in LQ games have focused on policy gradient approaches. While the convergence of policy gradient and its variants in single-agent infinite-horizon LQ Gaussian control problems is well-established [8], the case of LQ games has proven more challenging. In particular, in infinite-horizon general-sum LQ game settings,  [16] and [17] argued that there are neither global nor local convergence guarantees to a Nash equilibrium for policy gradient. This negative result has motivated several follow-up studies on understanding the convergence of policy gradient in LQ games.
For finite-horizon general-sum LQ games, the global convergence of the natural policy gradient method to the Nash equilibrium is proven in [10]. However, this result relies on the assumption of sufficient noise for exploration. One study showed that the policy gradient algorithm has convergence guarantees when agents interact through the aggregated state and action [23]. Other studies considered zero-sum [26], mean-field [5], or risk-sensitive [29] LQ games, proving the convergence of policy gradient variants in these settings.
In Markov games, of which LQ games are a subset, the challenges in proving convergence to Nash equilibria have motivated researchers to consider specific subclasses, particularly potential dynamic games. In static games, potential games are known for several desirable properties such as the existence of pure strategy Nash equilibria, the convergence of the best response dynamics [18], and learning dynamics to such equilibria [11]. These advances also have solid practical relevance in real-world dynamic games. The LQ game is a well-known dynamic game setting for many real-world applications, including congestion games [22], resource allocation games [14], games arising in multi-vehicle autonomous driving [1, 13], and electricity markets [6, 20]. Consequently, a natural approach to extend the convergence guarantees of learning-based methods to dynamic games is to characterize the set of dynamic games that admit a potential function.
Recent literature in the multi-agent learning community has focused on conditions that ensure a dynamic game with finite state and action spaces is potential. Studies on Markov games with finite states and actions [15, 27, 28] reveal that deriving conditions to characterize a potential Markov game is not straightforward. In particular, [15] demonstrated that the stage cost being potential is neither necessary nor sufficient for the dynamic game to be potential. Given these limitations, [27] focused on Markov games with decoupled dynamics and local policies where agents make decisions using only their local state. In this setting, they proved the convergence of a distributed learning algorithm to a Nash equilibrium for Markov potential games. By characterizing LQ potential games as a subset of Markov potential games, we gain a fundamental understanding of a significant class of games in control theory.
As a first step toward understanding Markov potential games in infinite state and action spaces, we consider LQ games. Existing results on LQ potential games are limited. In [25], a theoretical analysis of dynamic potential games with constrained state-action sets for LQ games is provided by extending the definition of static potential games. In [2], Bayesian optimization is used to develop a novel algorithm for computing the Nash equilibrium in LQ potential games. Both studies provided some examples of LQ potential games beyond identical interest games. However, in our current work, we challenge their characterization of potential games (see Remark 6) aiming to clarify a misunderstanding in previous works on the characterization of potential games.
1.1 Contributions
Generally, the objective of this paper is to advance a fundamental understanding of classes of LQ potential games. Our contributions towards this goal are as follows:
  • We derive conditions under which finite-horizon general-sum LQ games admit a potential function (see Lemma 4). Leveraging this result, we show that the LQ potential games with coupled dynamics and a full-state feedback information structure form a limiting class, meaning they are ”almost” the same as identical interest games (see Proposition 5 for precise statement).
  • Motivated by the aforementioned limitations, we consider a subclass of LQ games with agents having decoupled dynamics and decoupled information structure but coupled loss functions. For this subclass, we characterize the set of LQ potential games and show that this set is significantly larger than the set of identical interest games (see Theorem 9).
  • For potential games within this subclass, we derive the potential function (Proposition 11) and show the existence of Nash equilibria (Proposition 12). Furthermore, we formalize the challenges associated with proving the uniqueness of the Nash equilibrium and its computation theoretically (Proposition 13 and Section 4.2) and through simulation (Section 4.4).
The rest of the paper is organized as follows. In Section 2, we introduce the class of games under study and provide background information on potential games. In Section 3, we derive conditions for LQ potential games with coupled dynamics and full-state feedback and characterize the potential games in the case of decoupled dynamics and information structure. The properties of the LQ potential game, its corresponding structured optimal control problem, and the convergence of policy gradient for this control problem to a stationary point are provided in Section 4. We conclude and discuss open research directions in Section 5.
1.2 Notations
[…]
2 Linear quadratic game setup
We define a finite-horizon linear quadratic (LQ) game and a variant of it with agents having decoupled dynamics. The notion of Nash equilibrium under different information structures and the definition of LQ potential games are provided. Our first result, Lemma 4, derives conditions for an LQ game to be potential. This lemma serves as a foundation for deriving necessary and sufficient conditions for LQ potential games throughout this work.
2.1 Linear quadratic games
Consider a non-cooperative general-sum LQ game where the dynamics are defined as
$x_{t+1}=Ax_{t}+\sum_{i=1}^{N}B^{i}u^{i}_{t},\quad x_{0}\sim\mathcal{D},$ (1)
with $t\in\{0,\dots,t_{f}-1\}$. For simplicity, let us consider scalar actions denoted by $u^{i}_{t}\in\mathbb{R}$, and the states of the game are $x_{t}\in\mathbb{R}^{n}$. The sub-index and super-index indicate the time and the agent, respectively. The initial state $x_{0}\in\mathbb{R}^{n}$ is randomly distributed according to a distribution $\mathcal{D}$. We assume that all agents choose their decision variables simultaneously at each time step.
The action of agent $i$, $u^{i}_{t}$, depends on the information available to the agent. Let us denote the decision variable of agent $i$ as a function of its information by $\gamma^{i}_{t}\in\mathbb{R}^{q^{i}}$. The exact form of $\gamma^{i}$ will be detailed in Section 2.2. The loss function of each agent is
$J^{i}(\gamma)=$ $\mathop{\mathbb{E}}_{x_{0}\sim\mathcal{D}}\bigl[J^{i}_{x}(\gamma,x_{0})+J^{i}_{u}(\gamma,x_{0})\bigr],$ (2)
where the state part $J^{i}_{x}(\gamma,x_{0})\in\mathbb{R}$ and the action part $J^{i}_{u}(\gamma,x_{0})\in\mathbb{R}$ of the loss function are defined as follows:
$J^{i}_{x}(\gamma,x_{0}):=$ $\sum_{t=0}^{t_{f}}(x_{t}-d_{t})^{T}Q_{t}^{i}(x_{t}-d_{t}),$ (3)
$J^{i}_{u}(\gamma,x_{0}):=$ $\sum_{t=0}^{t_{f}-1}\sum_{j=1}^{N}\sum_{h=1}^{N}(u_{t}^{j})^{T}(R^{i}_{t})_{jh}u^{h}_{t},$ (4)
[…]
We say that the LQ game in (22) and (2) has decoupled dynamics if for $i\in[N]$ and $t\in\{0,\dots,t_{f}-1\}$,
$x_{t+1}^{i}=$ $A^{i}x^{i}_{t}+b^{i}u_{t}^{i},$ (5)
where $x_{t}^{i}\in\mathbb{R}^{n^{i}}$, $A^{i}\in\mathbb{R}^{n^{i}\times n^{i}}$, $b^{i}\in\mathbb{R}^{n^{i}}$ with $\sum_{i=1}^{N}n^{i}=n$. Note that the decoupled dynamics can be written as the dynamics introduced in (22) by defining the joint state as $x_{t}=(x_{t}^{1},\dots,x_{t}^{N})$, joint control as $u_{t}=(u_{t}^{1},\dots,u_{t}^{N})$, and $A=\text{blockdiag}\{A^{1},\dots,A^{N}\}\in\mathbb{R}^{n\times n}$, $B^{i}=(\mathbb{0}_{n^{1:i-1}},b^{i},\mathbb{0}_{n^{i+1:N}})\in\mathbb{R}^{n}$, for $i\in[N]$.
The motivation for considering LQ games with decoupled dynamics is to derive less restrictive conditions for the existence of a potential function. As demonstrated in the following two examples, this class of LQ games also holds practical relevance.
Example 1: In the context of decentralized formation control of multi-vehicle systems, each vehicle can be considered an agent, with $x_{t}^{i}$ and $u_{t}^{i}$ representing the position and control action of the $i$th vehicle at time step $t$. As noted in [1], one approach to defining each vehicle’s objective is to minimize the formation error and energy consumed by the vehicle. In [1], the formation error of the $i$th vehicle is defined as follows:
$e^{i}_{x}(x^{1}_{t},\dots,x^{N}_{t})\ =\sum_{j=1}^{N}w^{ij}_{t}\|x_{t}^{i}-x^{j}_{t}-d^{ij}_{t}\|^{2},$
where $d_{t}^{ij}=d^{i}_{t}-d^{j}_{t}$ indicates the desired distance between two vehicles. The error weights are $w^{ij}_{t}\in\mathbb{R}_{\geq 0}$. The energy consumption of vehicle $i$ at time $t$ is presented by the quadratic form of control actions $(u^{i}_{t})^{T}(R^{i}_{t})_{ii}u^{i}_{t}$ where $(R^{i}_{t})_{jh}$ is nonzero if and only if $j=i$ and $h=i$. The agents’ objective is the summation of formation error and energy consumption over a finite horizon and thus, can be written as in (2).
Example 2: Let us consider a simple dynamic Cournot game, where agent $i$ is a firm that decides how much product $u^{i}_{t}\in\mathbb{R}$ to sell at each time step, given the amount of product stored $x^{i}_{t}\in\mathbb{R}_{\geq 0}$. A simple abstraction of the agent’s dynamics is
$x^{i}_{t+1}=x^{i}_{t}-u^{i}_{t}.$
The firm’s goal is to maximize its profit which is a function of the price of the product. This price is a decreasing function of the total production of all agents. As commonly done in a Cournot model in [24][14], and [20], let us consider a linear price at time $t$ of $p_{t}(u_{t})=-\alpha_{t}\sum_{i=1}^{N}u^{i}_{t}$ where $\alpha_{t}\in\mathbb{R}_{>0}$. Maximizing the profit is equivalent to minimizing the following loss function:
$J^{i}(\gamma)=$ $\mathop{\mathbb{E}}_{x_{0}\sim\mathcal{D}}\bigl[Q_{t_{f}}^{i}(x^{i}_{t_{f}}-d_{t_{f}}^{i})^{2}-\sum_{t=0}^{t_{f}-1}p_{t}(u_{t})u^{i}_{t}\bigr].$ (6)
[…]
2.2 Nash equilibria and potential game definitions
For the LQ games defined in (22) and (2) we consider three different information structures: (i) open-loop, (ii) full-state linear feedback, and (iii) decoupled state linear feedback. In each case, for $i\in[N]$, the decision variables are
(i) actions $\gamma_{t}^{i}=u_{t}^{i}\in\mathbb{R}$ with $q^{i}=1$;
(ii) full-state linear feedback coefficients $\gamma_{t}^{i}=K^{i}_{t}\in\mathbb{R}^{n}$, resulting in $u_{t}^{i}=-K_{t}^{i}x_{t}$ and $q^{i}=n$;
(iii) decoupled state linear feedback coefficients $\gamma_{t}^{i}=k^{i}_{t}\in\mathbb{R}^{n^{i}}$, resulting in $u_{t}^{i}=-k_{t}^{i}x^{i}_{t}$ and $q^{i}=n^{i}$.
Definition 1.
The joint strategy $({\gamma}^{1*},\dots,{\gamma}^{N*})\in\mathbb{R}^{t_{f}q^{1:N}}$ is a Nash equilibrium (NE) for the LQ game defined in (22) and (2) if and only if for $i\in[N]$,
$J^{i}(\gamma^{i*},\gamma^{-i*})\leq J^{i}(\tilde{\gamma}^{i},\gamma^{-i*}),\qquad\forall\tilde{\gamma}^{i}\in\mathbb{R}^{t_{f}q^{i}}.$
Definition 2.
The LQ game defined in (22) and (2) is a potential game if and only if there exists a function $\Pi:\mathbb{R}^{t_{f}q^{1:N}}\rightarrow\mathbb{R}$ such that for $i\in[N]$, $\gamma^{-i}\in\mathbb{R}^{t_{f}\sum_{j\neq i}^{N}q^{j}}$, and $\hat{\gamma}^{i},{\gamma}^{i}\in\mathbb{R}^{t_{f}q^{i}}$,
$J^{i}(\gamma^{i},\gamma^{-i})-J^{i}(\hat{\gamma}^{i},\gamma^{-i})=\Pi(\gamma^{i},\gamma^{-i})-\Pi(\hat{\gamma}^{i},\gamma^{-i}).$ (7)
In this case, $\Pi$ is called a potential function for the game.
Identical interest LQ games defined below are trivial examples of potential games. While their definition is apparent from their name, to clarify in comparing our results we formalize them below.
Definition 3.
The LQ games in (22) and (2) are identical interest if for $i,j,h,l\in[N]$ and for $t\in\{0,\dots,t_{f}-1\}$,
$Q^{i}_{t+1}=Q^{j}_{t+1},\quad\text{and}\quad(R^{i}_{t})_{lh}=(R^{j}_{t})_{lh}.$ (8)
Our focus is on deriving conditions to identify non-trivial LQ potential games. Therefore, we derive the necessary and sufficient conditions for an LQ game to be potential.
3 LQ potential games
First, in Lemma 4, we develop a test to verify whether an LQ game with the three information structures above is potential. Using this lemma, we demonstrate that, unlike the open-loop information structure in (i), a game with full-state feedback information structure (ii) is potential if and only if agents’ losses are identical at each stage, except for the initial stage. This motivates us to consider decoupled dynamics and information structure (iii) defined in Subsection 3.3. For this subset of LQ games, we characterize conditions under which the game admits a potential function, beyond the identical interest loss function.
3.1 Necessary and sufficient conditions for LQ potential games
The game pseudo-gradient, $\mathcal{G}(\gamma):\mathbb{R}^{t_{f}q^{1:N}}\to\mathbb{R}^{t_{f}q^{1:N}}$, is defined as the gradient of each agent loss function with respect to her own decision variable as follows:
$\mathcal{G}(\gamma)=\begin{bmatrix}\big(\frac{\partial J^{1}(\gamma)}{\partial\gamma^{1}}\big)^{T},\dots,\big(\frac{\partial J^{N}(\gamma)}{\partial\gamma^{N}}\big)^{T}\end{bmatrix}^{T},$
where $\frac{\partial J^{i}(\gamma)}{\partial\gamma^{i}}\in\mathbb{R}^{t_{f}q^{i}}$. For a differentiable $\mathcal{G}(\gamma)$, the Jacobian of the pseudo-gradient, $\mathcal{J}:\mathbb{R}^{t_{f}q^{1:N}}\to\mathbb{R}^{t_{f}q^{1:N}\times t_{f}q^{1:N}}$, is:
$\mathcal{J}(\gamma)=\begin{bmatrix}\frac{\partial^{2}J^{1}(\gamma)}{\partial\gamma^{1}\partial\gamma^{1}}&\dots&\frac{\partial^{2}J^{1}(\gamma)}{\partial\gamma^{N}\partial\gamma^{1}}\\ \vdots&\ddots&\vdots\\ \frac{\partial^{2}J^{N}(\gamma)}{\partial\gamma^{1}\partial\gamma^{N}}&\dots&\frac{\partial^{2}J^{N}(\gamma)}{\partial\gamma^{N}\partial\gamma^{N}}\end{bmatrix}$, (9)
where $\frac{\partial^{2}J^{i}(\gamma)}{\partial\gamma^{j}\partial\gamma^{i}}\in\mathbb{R}^{t_{f}q^{i}\times t_{f}q^{j}}$. Our first result develops an equivalent condition to an LQ game being potential.
Lemma 4.
The LQ game defined in (22) and (2) is a potential game if and only if for ${\gamma}_{t}^{i}\in\mathbb{R}^{q^{i}}$, ${\gamma}_{\tau}^{j}\in\mathbb{R}^{q^{j}}$, $i,j\in[N]$, and $t,\tau\in\{0,\dots,t_{f}-1\}$
$\frac{\partial^{2}J^{i}(\gamma)}{\partial\gamma_{\tau}^{j}\partial\gamma_{t}^{i}}=\frac{\partial^{2}J^{j}(\gamma)}{\partial\gamma_{\tau}^{j}\partial\gamma_{t}^{i}}.$ (10)
{pf}
A static differentiable game is potential if and only if the Jacobian of its pseudo-gradient is symmetric [7, Theorem 1.3.1]. Our goal is now to verify that the condition in (10) is equivalent to the game Jacobian being symmetric. The block $ij$ of the matrix in (9) is:
$\frac{\partial^{2}J^{i}(\gamma)}{\partial\gamma^{j}\partial\gamma^{i}}=\begin{bmatrix}\frac{\partial^{2}J^{i}(\gamma)}{\partial\gamma_{0}^{j}\partial\gamma_{0}^{i}}&\dots&\frac{\partial^{2}J^{i}(\gamma)}{\partial\gamma_{t_{f}-1}^{j}\partial\gamma_{0}^{i}}\\ \vdots&\ddots&\vdots\\ \frac{\partial^{2}J^{i}(\gamma)}{\partial\gamma_{0}^{i}\partial\gamma_{t_{f}-1}^{i}}&\dots&\frac{\partial^{2}J^{i}(\gamma)}{\partial\gamma_{t_{f}-1}^{j}\partial\gamma_{t_{f}-1}^{i}}\end{bmatrix}$.
Since $J^{i}(\gamma)$ is a polynomial function of $\gamma$, it is twice-continuously differentiable, and thus
$\frac{\partial^{2}J^{i}(\gamma)}{\partial\gamma^{j}\partial\gamma^{i}}=\left[\frac{\partial^{2}J^{i}(\gamma)}{\partial\gamma^{i}\partial\gamma^{j}}\right]^{T}.$
Hence, the terms on the diagonal of the Jacobian in (9) are symmetric. Since diagonal terms are symmetric, the Jacobian is symmetric if and only if
$\frac{\partial^{2}J^{i}(\gamma)}{\partial\gamma^{j}\partial\gamma^{i}}=\left[\frac{\partial^{2}J^{j}(\gamma)}{\partial\gamma^{i}\partial\gamma^{j}}\right]^{T}.$
Note that the left-hand side of equation (10) is the entry in the $t$-th row and $\tau$-th column of $\frac{\partial^{2}J^{i}(\gamma)}{\partial\gamma^{j}\partial\gamma^{i}}$ and the right-hand side is the entry in the $t$-th row and $\tau$-th column of $\frac{\partial^{2}J^{j}(\gamma)}{\partial\gamma^{i}\partial\gamma^{j}}$, and $\frac{\partial^{2}J^{j}(\gamma)}{\partial\gamma^{i}\partial\gamma^{j}}$ is symmetric. Thus, the equality above holds if and only if conditions in (10) are met. ∎
Note that each element of $\mathcal{J}(\gamma)$ is a second derivative of the loss function $J^{i}(\gamma)$, and the gradients of the loss functions can be more than second-order polynomials with respect to the decision variable, depending on the information structure. Therefore, investigating conditions for desirable properties in the Jacobian, such as diagonally strictly concave [21], can be as challenging as deriving LQ potential conditions. Next, we focus on investigating the potential condition in Lemma 4 for different game settings and information structures.
3.2 Potential games with coupled dynamics
Lemma 4 provides a test for verifying whether an LQ is potential. However, computing the needed gradients in state-feedback information structure in general is challenging. Thus, we focus on a scalar state setting. In this setting, for full-state feedback information structure, we show that the class of potential games is almost the same as the class of identical interest games (see conditions in Proposition 5).
Example 3. Here we consider an LQ game defined by equations (22) and (2) where $N=2$, $n=1$, and $d_{t}^{1}=d_{t}^{2}=0$, and the dynamics is $x_{t+1}=ax_{t}+b^{1}u^{1}_{t}+b^{2}u^{2}_{t}$ for $t\in\{0,\dots,t_{f}-1\}$. In loss function in (2), for open-loop information structure (i), $\gamma_{t}^{i}=u_{t}^{i}$, and for full-state feedback (ii), $\gamma_{t}^{i}=K_{t}^{i}$ where $u_{t}^{i}=-K_{t}^{i}x_{t}$, for $i\in\{1,2\}$.
Proposition 5.
For the setting described in Example 3,
  • (a)
    the game is an open-loop potential game (information structure (i)) if and only if
    $Q^{1}_{1}b^{1}b^{2}+(R^{1}_{0})_{12}=Q^{2}_{1}b^{1}b^{2}+(R^{2}_{0})_{12}.$ (11)
    and for $t\in\{1,\dots,t_{f}-1\}$
    $Q^{1}_{t+1}=Q^{2}_{t+1},\quad\text{and}\quad(R_{t}^{1})_{12}=(R_{t}^{2})_{12}.$ (12)
  • (b)
    the game is potential with full-state feedback information structure (ii) if and only if for $t=0$, equation (11) holds, and for $t\in\{1,\dots,t_{f}-1\}$ and for $l,h\in\{1,2\}$,
    $Q^{1}_{t+1}=Q^{2}_{t+1},\quad\text{and}\quad(R^{1}_{t})_{lh}=(R^{2}_{t})_{lh}.$ (13)
Remark 6.
Let us clarify a misunderstanding in the past works on LQ potential games. Past works [2, 25] claimed that the LQ games in (2) with $Q_{t}^{i}=Q_{t}^{j}$, $(R^{i}_{t})_{jh}=0$ for $j\neq i$ and $h\neq i$, and any $(R^{i}_{t})_{ii}$ for $i\in[N]$, are LQ potential game under full-state feedback structure. However, our Proposition 5 demonstrates that such an LQ game is an open-loop potential but not potential with a full-state feedback structure. Specifically, the loss parameters of quadratic terms also need to be identical in the latter case, i.e., $(R^{i}_{t})_{ii}=(R^{j}_{t})_{ii}$ for $i,j\in[N]$. This condition is not met in their provided game setting since $(R^{j}_{t})_{ii}=0$ for $j\neq i$ and $(R^{i}_{t})_{ii}$ is not necessarily zero.
Before presenting the proof, some insight into the derived results is offered. Notably, for $t>0$, the stage losses must be identical for the information structure (ii) whereas for the information structure (i) the loss of agent $i$ on its control action need not be the same as that of another agent. The condition on stage loss parameters for the first step $t=0$ is different than for $t>0$ due to the varying dependency of stage losses on $\gamma_{t}$, for $t>0$. In particular, the first stage loss only depends on $\gamma_{0}$, and the last stage loss depends on $\gamma_{t}$, for $t\in\{0,\dots,t_{f}-1\}$. Consequently, the conditions on the parameters of the first stage loss appear only when evaluating the condition in (10) for $(t,\tau)=(0,0)$.
{pf}
For scalar states and actions, the loss function of agent $i$ in (2) where $i\in\{1,2\}$ for $t_{f}=2$ is
$J^{i}(\gamma)$ $=\mathop{\mathbb{E}}_{x_{0}\sim\mathcal{D}}\Big[\sum_{t=0}^{2}Q_{t}^{i}(x_{t})^{2}+\sum_{t=0}^{1}\Big((R_{t}^{i})_{11}(u_{t}^{1})^{2}$
$+2(R_{t}^{i})_{12}u_{t}^{1}u_{t}^{2}+(R_{t}^{i})_{22}(u_{t}^{2})^{2}\Big)\Big],$ (14)
where $x_{t+1}=a^{t}x_{0}+\sum_{\tau=0}^{t}a^{t-\tau}[b^{1}u^{1}_{\tau}+b^{2}u^{2}_{\tau}]$ for open-loop information structure (i), and $x_{t+1}=\prod_{\tau=0}^{t}(a-b^{1}K_{\tau}^{1}-b^{2}K^{2}_{\tau})x_{0}$ for full-state feedback (ii). By leveraging Lemma 4, the above decomposition is used in the proof.
(a) Open-loop information structure (i): First, we show that for $t_{f}=2$, the game is potential if and only if (11) is met for $t=0$, and conditions in (12) are satisfied for $t=1$. Subsequently, through induction, we demonstrate the validity of the proposition for any finite $t_{f}>2$.
For $t_{f}=2$, we verify that the condition in (10) is satisfied for the four pairs of time steps $(t,\tau)$ where $t,\tau\in\{0,1\}$, for $i=1$, and for $j=2$ if and only if conditions in Proposition 5 part (a) is held. In this proof, the derivatives of states are utilized, $\frac{\partial x_{1}}{\partial u_{0}^{i}}=b^{i}$, $\frac{\partial x_{1}}{\partial u_{1}^{i}}=0$, $\frac{\partial x_{2}}{\partial u_{0}^{i}}=ab^{i}$, and $\frac{\partial x_{2}}{\partial u_{1}^{i}}=b^{i}$ for $i\in\{1,2\}$. Observe that
$\frac{\partial J^{1}(\gamma)}{\partial u_{1}^{1}}$ $=2\mathop{\mathbb{E}}_{x_{0}\sim\mathcal{D}}\left[Q_{2}^{1}x_{2}\frac{\partial x_{2}}{\partial u_{1}^{1}}+(R^{1}_{1})_{11}u^{1}_{1}+(R^{1}_{1})_{12}u^{2}_{1}\right].$
The second derivatives of $J^{1}(\gamma)$ are
$\frac{\partial^{2}J^{1}(\gamma)}{\partial u_{1}^{2}\partial u_{1}^{1}}$ $=2\bigl(Q_{2}^{1}b^{1}b^{2}+(R_{1}^{1})_{12}\bigr),$
$\frac{\partial^{2}J^{1}(\gamma)}{\partial u_{0}^{2}\partial u_{1}^{1}}$ $=2Q_{2}^{1}ab^{1}b^{2}.$
If $(R_{1}^{1})_{12}$ and $Q_{2}^{1}$ are respectively replaced by $(R_{1}^{2})_{12}$ and $Q_{2}^{2}$ in the equations above, the second derivatives of $J^{2}(\gamma)$ with respect to $u^{1}_{1}$ and $u^{2}_{\tau}$, for $\tau\in\{0,1\}$ are achieved. Condition in (10) is met for two pairs of $(t,\tau)\in\{(1,1),(1,0)\}$ if and only if $Q_{2}^{1}ab^{1}b^{2}=Q_{2}^{2}ab^{1}b^{2}$, and $Q_{2}^{1}b^{1}b^{2}+(R_{1}^{1})_{12}=Q_{2}^{2}b^{1}b^{2}+(R_{1}^{2})_{12}$. Since states are scalar, and dynamics parameters $b^{1}$ and $b^{2}$ are nonzero, these two conditions are met if and only if $Q_{2}^{1}=Q_{2}^{2}$ and $(R_{1}^{1})_{12}=(R_{1}^{2})_{12}$ that are equivalent to (12) for $t=1$.
The derivative of $J^{1}(\gamma)$ with respect to $u_{0}^{1}$ is
$\frac{\partial J^{1}(\gamma)}{\partial u_{0}^{1}}$ $=2\mathop{\mathbb{E}}_{x_{0}\sim\mathcal{D}}\biggl[Q_{2}^{1}x_{2}\frac{\partial x_{2}}{\partial u_{0}^{1}}+Q_{1}^{1}x_{1}\frac{\partial x_{1}}{\partial u_{0}^{1}}$
$+(R^{1}_{0})_{11}u^{1}_{0}+(R^{1}_{0})_{12}u^{2}_{0}\biggr].$
One can simply check that for the pair of $(t,\tau)=(0,1)$ in (10), the repetitive condition $Q_{2}^{1}=Q_{2}^{2}$ will be derived. The second derivative of $J^{1}(\gamma)$ with respect to $u_{0}^{2}$ is
$\frac{\partial^{2}J^{1}(\gamma)}{\partial u_{0}^{2}\partial u_{0}^{1}}$ $=2\bigl[(a)^{2}Q^{1}_{2}b^{1}b^{2}+Q^{1}_{1}b^{1}b^{2}+(R^{1}_{0})_{12}\big].$
If $(R_{0}^{1})_{12}$, $Q_{2}^{1}$, and $Q_{1}^{1}$ are respectively replaced by $(R_{0}^{2})_{12}$, $Q_{2}^{2}$, and $Q_{1}^{2}$ in the equation above, the second derivatives of $J^{2}(\gamma)$ with respect to $u^{1}_{0}$ and $u^{2}_{0}$ is achieved. Since we already established that $Q_{2}^{1}=Q_{2}^{2}$, for $(t,\tau)=(0,0)$, the condition in (10) is satisfied if and only if
$Q^{1}_{1}b^{1}b^{2}+(R^{1}_{0})_{12}=Q^{2}_{1}b^{1}b^{2}+(R^{2}_{0})_{12},$
where is identical to (11). By consolidating all necessary and sufficient conditions for different pairs of $(t,\tau)$, the Proposition 5 part (a) is proved for $t_{f}=2$.
The proof can be extended for $t_{f}>2$ through induction. We assume that conditions in Proposition 5 part (a) are necessary and sufficient conditions for the LQ game in Example 2 with $t_{f}=h$ with open-loop information structure to be potential, and we aim to demonstrate that these conditions are also necessary and sufficient conditions for the game with $t_{f}=h+1$. The difference between the loss function of agent 1 for $t_{f}=h$ and for $t_{f}=h+1$ lies in the last stage loss $Q_{h+1}^{1}(x_{h+1})^{2}+(R_{h}^{1})_{11}(u_{h}^{1})^{2}+2(R_{h}^{1})_{12}u_{h}^{1}u_{h}^{2}+(R_{h}^{1})_{22}(u_{h}^{2})^{2}$. The decision variables $u_{h}^{1}$ and $u_{h}^{2}$ only appear in the last stage loss so that the second derivative of $J^{1}(\gamma)$ is
$\frac{\partial^{2}J^{1}(\gamma)}{\partial u_{h}^{2}\partial u_{h}^{1}}$ $=2\big[Q_{h+1}^{1}b^{1}b^{2}+(R_{h}^{1})_{12}\big],$
$\frac{\partial^{2}J^{1}(\gamma)}{\partial u_{h-1}^{2}\partial u_{h}^{1}}$ $=2Q_{h+1}^{1}ab^{1}b^{2}.$
Therefore, for the pair of $(t,\tau)\in\{(h,h),(h,h-1)\}$, the condition in (10) is met if and only if $Q_{h+1}^{1}=Q_{h+1}^{2}$ and $(R_{h}^{1})_{12}=(R_{h}^{2})_{12}$. The contribution of the last stage loss in the conditions obtained by other pairs of $(t,\tau)$ is repetitive, and other stage losses result in similar conditions derived for the case where $t_{f}=h$. Therefore, the game described in Example 3 with $t_{f}=h+1$ is open-loop potential if and only if conditions in Proposition 5 part (a) are satisfied.
(b) Full-state feedback information structure (ii): First, we prove that for $t_{f}=2$, the game is potential if and only if equation (11) is met for $t=0$, and conditions in (13) are satisfied for $t=1$. Then we utilize induction to demonstrate that Proposition 5 part (b) is also true for $t_{f}>2$.
For $t_{f}=2$, we examine that the condition in (10) is satisfied for the four pairs of time steps $(t,\tau)$ where $t,\tau\in\{0,1\}$ for $i=1$, and for $j=2$ if and only if conditions in Proposition 5 part (b) are held. In this proof, the derivatives of states, $\frac{\partial x_{1}}{\partial K_{0}^{i}}=-b^{i}x_{0}$, $\frac{\partial x_{1}}{\partial K_{1}^{i}}=0$, $\frac{\partial x_{2}}{\partial K_{0}^{i}}=-b^{i}(a-b^{1}K_{1}^{1}-b^{2}K_{1}^{2})x_{0}$, and $\frac{\partial x_{2}}{\partial K_{1}^{i}}=-b^{i}x_{1}$, and nonzero derivatives of actions, $\frac{\partial u^{i}_{1}}{\partial K_{0}^{j}}=K^{i}_{1}b^{j}x_{0}$, $\frac{\partial u^{i}_{1}}{\partial K_{1}^{i}}=-x_{1}$, and $\frac{\partial u^{i}_{0}}{\partial K_{0}^{i}}=-x_{0}$, are utilized for $i,j\in\{1,2\}$. Here, we only derive $J^{1}(\gamma)$ derivatives, but similar to the open-loop case, the $J^{2}(\gamma)$ derivative can be computed by replacing the super-index 1 with 2 for $Q$ and $R$ parameters. The derivative of $J^{1}(\gamma)$ in (14) with respect to $K_{1}^{1}$ is
$\frac{\partial J^{1}(\gamma)}{\partial K_{1}^{1}}$ $=2\mathop{\mathbb{E}}_{x_{0}\sim\mathcal{D}}\biggl[Q_{2}^{1}x_{2}\frac{\partial x_{2}}{\partial K_{1}^{1}}$
$+\{(R^{1}_{1})_{11}u^{1}_{1}+(R^{1}_{1})_{12}u^{2}_{1}\}\frac{\partial u^{1}_{1}}{\partial K_{1}^{1}}\biggr].$
The derivative of $\frac{\partial J^{1}(\gamma)}{\partial K_{1}^{1}}$ with respect to $K_{1}^{2}$ and $K_{0}^{2}$ are
$\frac{\partial^{2}J^{1}(\gamma)}{\partial K_{1}^{2}\partial K_{1}^{1}}$ $=2[Q_{2}^{1}b^{1}b^{2}+(R_{1}^{1})_{12}]\mathop{\mathbb{E}}_{x_{0}\sim\mathcal{D}}\left[(x_{1})^{2}\right],$
$\frac{\partial^{2}J^{1}(\gamma)}{\partial K_{0}^{2}\partial K_{1}^{1}}$ $=4\big[Q_{2}^{1}b^{1}b^{2}(a-b^{1}K_{1}^{1}-b^{1}K_{1}^{2})$
$-b^{2}\{(R_{1}^{1})_{11}K_{1}^{1}+(R_{1}^{1})_{12}K_{1}^{2}\}\big]\mathop{\mathbb{E}}_{x_{0}\sim\mathcal{D}}[x_{1}x_{0}].$
For $(t,\tau)=(1,0)$, the condition in (10) is met if and only if the coefficient of $K_{1}^{1}$ and $K_{1}^{2}$ and the constant term of $J^{1}(\gamma)$ and $J^{2}(\gamma)$ second derivatives are equal as follows.
$Q_{2}^{1}(b^{1})^{2}b^{2}$ $=Q_{2}^{2}(b^{1})^{2}b^{2}$
$Q_{2}^{1}(b^{1})^{2}b^{2}+b^{2}(R_{1}^{1})_{11}$ $=Q_{2}^{2}(b^{1})^{2}b^{2}+b^{2}(R_{1}^{2})_{11}$
$Q_{2}^{1}b^{1}(b^{2})^{2}+b^{2}(R_{1}^{1})_{12}$ $=Q_{2}^{2}b^{1}(b^{2})^{2}+b^{2}(R_{1}^{2})_{12}.$
Thus, the condition in (10) is met for $(t,\tau)=(1,0)$ if and only if $Q_{2}^{1}=Q_{2}^{2}$, $(R_{1}^{1})_{12}=(R_{1}^{2})_{12}$, and $(R_{1}^{1})_{11}=(R_{1}^{2})_{11}$, which are comparable with (13) for $t=1$. Consequently, the second derivatives of $J^{1}(\gamma)$ and $J^{2}(\gamma)$ with respect to $K_{1}^{1}$ and $K_{1}^{2}$ are also equal.
The first derivative of $J^{1}(\gamma)$ in (14) with respect to $K_{0}^{1}$ is
$\frac{\partial J^{1}(\gamma)}{\partial K_{0}^{1}}$ $=2\mathop{\mathbb{E}}_{x_{0}\sim\mathcal{D}}\biggl[Q_{2}^{1}x_{2}\frac{\partial x_{2}}{\partial K_{0}^{1}}+Q_{1}^{1}x_{1}\frac{\partial x_{1}}{\partial K_{0}^{1}}$
$+\{(R^{1}_{0})_{11}u^{1}_{0}+(R^{1}_{0})_{12}u^{2}_{0}\}\frac{\partial u^{1}_{0}}{\partial K_{0}^{1}}$
$+\{(R^{1}_{1})_{11}u^{1}_{1}+(R^{1}_{1})_{12}u^{2}_{1}\}\frac{\partial u^{1}_{1}}{\partial K_{0}^{1}}$
$+\{(R^{1}_{1})_{12}u^{1}_{1}+(R^{1}_{1})_{22}u^{2}_{1}\}\frac{\partial u^{2}_{1}}{\partial K_{0}^{1}}\biggr].$
The derivative of $\frac{\partial J^{1}(\gamma)}{\partial K_{0}^{1}}$ with respect to $K_{1}^{2}$ and $K_{0}^{2}$ are
$\frac{\partial^{2}J^{1}(\gamma)}{\partial K_{1}^{2}\partial K_{0}^{1}}$ $=4\big[Q_{2}^{1}b^{1}b^{2}(a-b^{1}K_{1}^{1}-b^{2}K_{1}^{2})$
$-b^{1}\{(R_{1}^{1})_{12}K_{1}^{1}+(R_{1}^{1})_{22}K^{2}_{1}\}\big]\mathop{\mathbb{E}}_{x_{0}\sim\mathcal{D}}[x_{1}x_{0}]$
$\frac{\partial^{2}J^{1}(\gamma)}{\partial K_{0}^{2}\partial K_{0}^{1}}$ $=2\{b^{1}b^{2}\big[Q_{1}^{1}+Q_{2}^{1}(a-b^{1}K_{1}^{1}-b^{2}K_{1}^{2})^{2}$
$+(R_{1}^{1})_{11}(K_{1}^{1})^{2}+2(R_{1}^{1})_{12}K^{1}_{1}K^{2}_{1}$
$+(R_{1}^{1})_{22}(K_{1}^{2})^{2}\big]+(R_{0}^{1})_{12}\}\mathop{\mathbb{E}}_{x_{0}\sim\mathcal{D}}[(x_{0})^{2}].$
Since we showed that $Q_{2}^{1}=Q_{2}^{2}$, $(R_{1}^{1})_{12}=(R_{1}^{2})_{12}$, and $(R_{1}^{1})_{11}=(R_{1}^{2})_{11}$, the condition in (10) is met for $(t,\tau)=(0,1)$ if and only if $(R_{1}^{1})_{22}=(R_{1}^{2})_{22}$. The equality in (10) is also true for $(t,\tau)=(0,0)$ if and only if
$Q^{1}_{1}b^{1}b^{2}+(R^{1}_{0})_{12}=Q^{2}_{1}b^{1}b^{2}+(R^{2}_{0})_{12},$
which is equivalent to Equation (11). Consolidating conditions derived for all pairs of $(t,\tau)$, we notice the game is potential with full-state feedback information structure if and only if equation (11) is met for $t=0$, and conditions in (13) are met for $t\in\{1,2\}$.
The proof can be extended for $t_{f}>2$ using induction. We assume that conditions in Proposition 5 part (b) are necessary and sufficient conditions for the LQ game in Example 2 with $t_{f}=h$ with full-state feedback information structure to be potential. Then, we aim to demonstrate that these conditions are necessary and sufficient for the game with $t_{f}=h+1$. The difference between $J^{1}(\gamma)$ for $t_{f}=h$ and $t_{f}=h+1$ is the last stage loss $Q_{h+1}^{1}(x_{h+1})^{2}+(R_{h}^{1})_{11}(u_{h}^{1})^{2}+2(R_{h}^{1})_{12}u_{h}^{1}u_{h}^{2}+(R_{h}^{1})_{22}(u_{h}^{2})^{2}$. One can simply check that
$\frac{\partial^{2}J^{1}(\gamma)}{\partial K_{h-1}^{2}\partial K_{h}^{1}}$ $=4\big[Q_{h+1}^{1}ab^{1}b^{2}$
$-\{Q_{h+1}^{1}(b^{1})^{2}b^{2}+b^{2}(R_{h}^{1})_{11}\}K_{h}^{1}$
$-\{Q_{h+1}^{1}b^{1}(b^{2})^{2}+b^{2}(R_{h}^{1})_{12}\}K_{h-1}^{2}\big]$
$\times\mathop{\mathbb{E}}_{x_{0}\sim\mathcal{D}}[x_{h-1}x_{h-2}],$
$\frac{\partial^{2}J^{1}(\gamma)}{\partial K_{h}^{2}\partial K_{h-1}^{1}}$ $=4\big[Q_{h+1}^{1}b^{1}b^{2}(a-b^{1}K_{h}^{1}-b^{2}K_{h}^{2})-b^{1}$
$\{(R_{h}^{1})_{12}K_{h}^{1}+(R_{h}^{1})_{22}K^{2}_{h}\}\big]\mathop{\mathbb{E}}_{x_{0}\sim\mathcal{D}}[x_{h}x_{h-1}].$
Therefore, for the pair of $(t,\tau)=(h,h-1)\}$, the condition in (10) is met if and only if $Q_{h+1}^{1}=Q_{h+1}^{2}$, $(R_{h}^{1})_{11}=(R_{h}^{2})_{11}$, $(R_{h}^{1})_{22}=(R_{h}^{2})_{22}$, and $(R_{h}^{1})_{12}=(R_{h}^{2})_{12}$. ∎
Remark 7.
We expect that extending the result to higher dimensions will not, in general, relax the conditions derived for being a potential game. However, extending the proof beyond the scalar case is challenging, if not very tedious, due to reasoning about the set of equalities that arise. Let us consider $t_{f}=2$. Observe that for higher dimensions of states, a set of dynamics parameters might exist such that the condition in (10) is met for this pair of $(t,\tau)$ even though $Q_{2}^{1}\neq Q_{2}^{2}$. Such cases might happen when $B^{i}(B^{i})^{T}$ is not full rank, and consequently, the following condition derived through (10) for the pair of $(t,\tau)=(1,0)$ may not have a unique solution.
$(B^{2})^{T}A^{T}Q_{2}^{1}B^{1}=(B^{2})^{T}A^{T}Q_{2}^{2}B^{1}.$
In summary, in higher dimensions, it might be possible to derive slightly relaxed, but system parameter-dependent ($A,B^{i}$) conditions equivalent to those in Proposition 5.
Remark 8.
To investigate the effect of the noisy dynamics on the results of Proposition 5, we add a random noise $w_{t}$ drawn from the distribution $\mathcal{W}$ with zero mean to the game dynamics in equation (22). The loss function will be defined as follows.
$J^{i}(\gamma)=$ $\mathop{\mathbb{E}}_{x_{0}\sim\mathcal{D}}\bigl[\sum_{t=0}^{t_{f}-1}\mathop{\mathbb{E}}_{w_{t}\sim\mathcal{W}}c_{t}(x_{t+1},u_{t})\bigr],$ (15)
where the stage loss $c_{t}(x_{t+1},u_{t})$ is defined the same as the one for the game without noise. Considering the loss function defined above, the definition of Nash equilibrium, potential game, identical interest games for LQ games (Definitions 1, 2, and 3), the game pseudo-gradient, $\mathcal{G}(\gamma)$, and its Jacobian in equation (9) do not change. Hence, Lemma 4 is still valid for the game with noisy dynamics. Note that the derivative of $c_{t}(x_{t+1},u_{t})$ with respect to $x_{t}$ or $u_{t}^{j}$ is similar to the game without noise. For open-loop information structure, one can show that the derivatives of $x_{t+1}$ and $u_{t}^{i}$ with respect to the decision variable $u_{\tau}^{j}$ do not depend on $w_{t}$, which indicates that the results of Proposition 5, part (a) is the same for games with and without noise. For full-state linear feedback information structure, one can show that the derivatives of $x_{t+1}$ and $u_{t}^{i}$ with respect to the decision variable $K_{\tau}^{j}$ are zero for $t+1\geq\tau$ and for $t+1<\tau$ can be written as a product of a coefficient that do not depend on $w_{t}$ and $x_{\tau}$. This observation implies that by investigating conditions in Lemma 4, the potential conditions derived for the LQ game with noisy dynamics and full-state feedback information structure are the same as those for the game without noise in Proposition 5, part (b). Details are provided in Appendix A.
The challenge of defining a class of potential games beyond identical interest games, coupled with the desirable properties of potential games (as discussed in the introduction), motivates us to restrict the game setting. Given the properties of Markov potential games with decoupled dynamics and local policies in finite state and action spaces [27], we confine the LQ game setting, where the state and action spaces are infinite, to decoupled dynamics and the decoupled state linear feedback information structure. Moreover, restricting the $A$ and $B^{i}$ matrices in LQ games might lead to a potential game that satisfies the conditions in Remark 7 without being an identical interest game. As described in Section 2, LQ games with decoupled dynamics are a practically relevant subset of LQ games. On the other hand, the closed-loop Nash equilibrium of the general-sum LQ game, and consequently the LQ game with decoupled dynamics, has a linear full-state feedback structure [4]. To relax the conditions described in Proposition 5, an information structure sparser than full-state linear feedback, such as decoupled state linear feedback, should be investigated. Thus, we focus on LQ games with decoupled dynamics and a decoupled state linear feedback information structure, similar to the study [27] on Markov potential games with decoupled dynamics and localized policies.
3.3 Decoupled dynamics and information structure
Here, we provide sufficient conditions for the LQ game with decoupled dynamics in (5) and decoupled information structure (iii) to be potential. Furthermore, we show that there exist practically relevant classes of LQ games that satisfy these conditions.
Theorem 9.
The LQ game defined in (2) with decoupled dynamics in (5) and the decoupled state feedback information structure in (iii) is potential if the following two conditions are satisfied for $i,j\in[N]$ and $t\in\{0,\dots,t_{f}-1\}$.
  1. (C1)
    $(Q^{i}_{t+1})_{ij}=(Q^{j}_{t+1})_{ij}:=Q_{t+1}^{ij}$, where $(Q^{i}_{t+1})_{hl}\in\mathbb{R}^{n^{h}\times n^{l}}$ is a sub-matrix of $Q_{t+1}^{i}$ consisting of the entries relevant to $i$ and $j$ agents as follows
    $Q^{i}_{t+1}=\begin{bmatrix}(Q^{i}_{t+1})_{11}&\dots&(Q^{i}_{t+1})_{1N}\\ \vdots&\ddots&\vdots\\ (Q^{i}_{t+1})_{N1}&\dots&(Q^{i}_{t+1})_{NN}\end{bmatrix}$. (16)
  2. (C2)
    $(R_{t}^{i})_{ij}=(R_{t}^{j})_{ij}:=R_{t}^{ij}$.
For a clear comparison of the potential conditions in Proposition 5 and Theorem 9, let us write the action part of the loss function in equation (4) in a closed form, $J^{i}_{u}(\gamma,x_{0})=\sum_{t=0}^{t_{f}-1}(u_{t})^{T}R^{i}_{t}u_{t}$ where
$R^{i}_{t}:=\begin{bmatrix}(R^{i}_{t})_{11}&\dots&(R^{i}_{t})_{1N}\\ \vdots&\ddots&\vdots\\ (R^{i}_{t})_{N1}&\dots&(R^{i}_{t})_{NN}\end{bmatrix}$, (17)
and $u=(u^{1},\dots,u^{N})\in\mathbb{R}^{N}$. Note that the state part of the loss function $J^{i}_{x}(\gamma,x_{0})$ in equation (3) can also be written in a closed form when the dynamics are decoupled using the definition of $Q_{t}^{i}$ in equation (16). Proposition 5 states that the game is potential with a full-state feedback information structure if and only if all loss function parameters for both players are the same (except for those corresponding to $t=0$), i.e., $R^{i}_{t}=R^{j}_{t}$ and $Q^{i}_{t}=Q^{j}_{t}$. However, Theorem 9 indicates that the game is potential with a decoupled state feedback information structure if the non-diagonal elements of $R^{i}_{t}$ and $Q^{i}_{t}$ are equal to the non-diagonal ones of $R^{j}_{t}$ and $Q^{j}_{t}$, respectively. It does not impose any conditions on the diagonal elements. As mentioned in Remark 6, the LQ games defined in [2, 25] with non-zero $(R^{i}_{t})_{ii}$ for $i\in[N]$ are assumed to be potential under full-state linear feedback structure if $Q_{t}^{i}=Q_{t}^{j}$ and non-diagonal elements of $R^{i}_{t}$ and $R^{j}_{t}$ are zero. Although the discussion in Remark 6 indicates that this claim is wrong, according to Theorem 9 and assuming decoupled dynamics, one can show that these games are an LQ potential game with decoupled state linear feedback information structure.
{pf}
The proof leverages Lemma 4. In particular, we verify that an LQ game with decoupled dynamics that satisfies conditions (C1) and (C2) will satisfy the conditions in (10), and therefore, is potential. Then, by providing more details on conditions, we show that (C1) and (C2) are not necessary for the games to be potential.
For the decoupled state feedback information structure in (iii), the decision variable is $\gamma_{t}^{i}=k_{t}^{i}$ where $u_{t}^{i}=-k^{i}_{t}x^{i}_{t}$. Since the dynamics are decoupled as equation (5), $u_{t}^{i}$ is the function of $(k^{i}_{0},\dots,k^{i}_{t})$; $x_{t}^{i}$ is the function of $(k^{i}_{0},\dots,k^{i}_{t-1})$; and both are independent of the other agents’ decision variable, $k^{-i}$. Consequently, for $i\in[N]$, $J_{x}^{i}(\gamma,x_{0})$ can be written as follows:
$J_{x}^{i}(\gamma,x_{0})$ $=\sum_{t=0}^{t_{f}}\sum_{j=1}^{N}\sum_{h=1}^{N}\big[(\bar{x}^{j}_{t})^{T}(Q^{i}_{t})_{jh}\bar{x}^{h}_{t}\big],$
where $\bar{x}^{j}_{t}=x^{j}_{t}-d^{j}_{t}$. For the state part of the loss function, the first derivative of $J^{i}_{x}(\gamma,x_{0})$ with respect to $k^{i}_{t}$ for $t\in\{0,\dots,t_{f}-1\}$ is computed using the chain rule and product rule as follow,
$\frac{\partial J_{x}^{i}(\gamma,x_{0})}{\partial k^{i}_{t}}=$ $\sum_{t^{\prime}=t+1}^{t_{f}}\sum_{h=1}^{N}2(\bar{x}_{t^{\prime}}^{h})^{T}(Q^{i}_{t^{\prime}})_{hi}\frac{\partial\bar{x}_{t^{\prime}}^{i}}{\partial k^{i}_{t}}.$
[…]
$\frac{\partial^{2}J_{x}^{i}(\gamma,x_{0})}{\partial k_{\tau}^{j}\partial k_{t}^{i}}=$ $\sum_{t^{\prime}=\text{max}(t,\tau)+1}^{t_{f}}2(\frac{\partial\bar{x}^{i}_{t^{\prime}}}{\partial k_{t}^{i}})^{T}(Q_{t^{\prime}}^{i})_{ji}^{T}\frac{\partial\bar{x}_{t^{\prime}}^{j}}{\partial k_{\tau}^{j}}.$
For the action part of the loss function in (4), similar steps are taken to compute the second derivatives of $J_{u}^{i}(\gamma,x_{0})$ with respect to $k^{i}_{t}$ and $k^{j}_{\tau}$ as follows
$\frac{\partial^{2}J_{u}^{i}(\gamma,x_{0})}{\partial k_{\tau}^{j}\partial k_{t}^{i}}=$ $\sum_{t^{\prime}=\text{max}(t,\tau)}^{t_{f}-1}2(R_{t^{\prime}}^{i})_{ji}(\frac{\partial u_{t^{\prime}}^{i}}{\partial k_{t}^{i}})^{T}\frac{\partial u^{j}_{t^{\prime}}}{\partial k_{\tau}^{j}},$
[…]
$\frac{\partial^{2}J^{i}(\gamma)}{\partial k_{\tau}^{j}\partial k_{t}^{i}}=$ $2\mathop{\mathbb{E}}_{x_{0}\sim\mathcal{D}}\Bigg[\sum_{t^{\prime}=\text{max}(t,\tau)}^{t_{f}-1}\{(\frac{\partial\bar{x}^{i}_{t^{\prime}+1}}{\partial k_{t}^{i}})^{T}(Q_{t^{\prime}+1}^{i})_{ji}^{T}$
$\frac{\partial\bar{x}_{t^{\prime}+1}^{j}}{\partial k_{\tau}^{j}}+(R_{t^{\prime}}^{i})_{ji}(\frac{\partial u_{t^{\prime}}^{i}}{\partial k_{t}^{i}})^{T}\frac{\partial u^{j}_{t^{\prime}}}{\partial k_{\tau}^{j}}\}\Bigg],$
If $(R_{t^{\prime}}^{i})_{ji}$ and $(Q_{t^{\prime}+1}^{i})_{ji}$ are respectively replaced by $(R_{t^{\prime}}^{j})_{ji}$ and $(Q_{t^{\prime}+1}^{j})_{ji}$ in the equation above, the second derivatives of $J^{j}(\gamma)$ with respect to $k^{i}_{t}$ and $k^{j}_{\tau}$ is achieved. Since the game setting satisfies (C1) and (C2), these loss parameters are equal, i.e. $(R_{t^{\prime}}^{i})_{ji}=(R_{t^{\prime}}^{j})_{ji}$ and $(Q_{t^{\prime}+1}^{i})_{ji}=(Q_{t^{\prime}+1}^{j})_{ji}$, for $t^{\prime}\in\{0,\dots,t_{f}-1\}$ and for $i,j\in[N]$. Accordingly, the second derivatives of $J^{j}(\gamma)$ and $J^{i}(\gamma)$ are equal, and consequently, the LQ game is potential.
Next, we show that conditions (C1) and (C2) are not necessary conditions for the game to be potential. For the pair $(t,\tau)=(t_{f}-1,t_{f}-1)$, the second derivative of $J^{i}(\gamma)$ with respect to $k^{i}_{t}$ and $k^{j}_{\tau}$ is
$\frac{\partial^{2}J^{i}(\gamma)}{\partial k_{t_{f}-1}^{j}\partial k_{t_{f}-1}^{i}}=$ $2\big\{(b^{i})^{T}(Q_{t_{f}}^{i})_{ij}b^{j}+(R_{t_{f}-1}^{i})_{ij}\big\}$
$\times\mathop{\mathbb{E}}_{x_{0}\sim\mathcal{D}}[x_{t_{f}-1}^{i}\big(x_{t_{f}-1}^{j}\big)^{T}],$
since the derivatives of state and action of agent $i$ are $\frac{\partial\bar{x}^{i}_{t_{f}}}{\partial k_{t_{f}-1}^{i}}=-b^{i}(x^{i}_{t_{f}-1})^{T}$ and $\frac{\partial{u}^{i}_{t_{f}-1}}{\partial k_{t_{f}-1}^{i}}=-(x^{i}_{t_{f}-1})^{T}$. For $(t,\tau)=(t_{f}-1,t_{f}-2)$, the second derivative of $J^{i}(\gamma)$ with respect to $k^{i}_{t}$ and $k^{j}_{\tau}$ is
$\frac{\partial^{2}J^{i}(\gamma)}{\partial k_{t_{f}-2}^{j}\partial k_{t_{f}-1}^{i}}=2\big\{(b^{i})^{T}(Q_{t_{f}}^{i})_{ij}A^{j}b^{j}-\big[(b^{i})^{T}(Q_{t_{f}}^{i})_{ij}b^{j}$
$+(R_{t_{f}-1}^{i})_{ij}\big]k_{t_{f}-1}^{j}b^{j}\big\}\mathop{\mathbb{E}}[x_{t_{f}-1}^{i}\big(x_{t_{f}-1}^{j}\big)^{T}],$
since the derivatives of states and action of agent $j$ are $\frac{\partial\bar{x}^{j}_{t_{f}}}{\partial k_{t_{f}-2}^{j}}=-(A^{j}-b^{j}k^{j}_{t_{f}-1})b^{j}(x^{j}_{t_{f}-2})^{T}$ and $\frac{\partial{u}^{j}_{t_{f}-1}}{\partial k_{t_{f}-2}^{j}}=k_{t_{f}-1}b^{j}(x^{j}_{t_{f}-2})^{T}$. From the pair of time steps $(t_{f}-1,t_{f}-1)$, the condition in (10) is met if and only if
$(b^{i})^{T}(Q_{t_{f}}^{i})_{ij}A^{j}b^{j}=$ $(b^{i})^{T}(Q_{t_{f}}^{j})_{ij}A^{j}b^{j},$
$(b^{i})^{T}(Q_{t_{f}}^{i})_{ij}b^{j}+(R_{t_{f}-1}^{i})_{ij}=$ $(b^{i})^{T}(Q_{t_{f}}^{j})_{ij}b^{j}+(R_{t_{f}-1}^{j})_{ij}.$
As discussed in Remark 7, this set of equations may not have a unique solution for $x^{i}$ with a dimension higher than one. Note that the term $(b^{i})^{T}((A^{i})^{t^{\prime}})^{T}(Q_{t_{f}}^{i})_{ij}(A^{j})^{t"}b^{j}$ where $t^{\prime},t"\in\{0,\dots,t_{f}-2\}$ will appear in several equations if other pairs of $(t,\tau)$ are calculated. This notion indicates that conditions on $A^{i}$ and $b^{i}$ might be very restrictive such that the LQ game with decoupled dynamics and decoupled feedback information structure satisfying conditions (C1) and (C2) might be the only game that meets condition (10) for all pairs of $(t,\tau)$.∎
Remark 10.
To investigate the effect of the noisy dynamics on the results of Theorem 9, we add a random noise $w^{i}_{t}$ drawn from the distribution $\mathcal{W}^{i}$ with zero mean to the set of decoupled dynamics in equation (5). By defining $w_{t}:=(w^{1}_{t},\dots,w^{N}_{t})$, the loss function of the LQ game with noisy decoupled dynamics can be written as the one in equation 15. Thus, Lemma 4 is still valid for the game with noisy dynamics and decoupled state linear feedback information structure, and the derivative of $c_{t}(x_{t+1},u_{t})$ with respect to $x_{t}$ or $u_{t}^{j}$ is similar to the game without noise. For decoupled state linear feedback information structure, the derivatives of $x^{i}_{t+1}$ and $u_{t}^{i}$ with respect to the decision variable $k_{\tau}^{i}$ are zero for $t+1\geq\tau$ and for $t+1<\tau$ can be written as a product of a coefficient that do not depend on $w_{t}$ and $x^{i}_{\tau}$. Note that the derivatives of $x^{i}_{t+1}$ and $u_{t}^{i}$ with respect to $k_{t}^{j}$ for $j\neq i$ are zero. This observation implies that by investigating conditions in Lemma 4, the potential conditions derived for the LQ game with noisy decoupled dynamics and full-state feedback information structure are the same as those for the game without noise in Theorem 9. Details are provided in Appendix A.
After theoretically deriving non-trivial (non-identical interest) classes of state-feedback potential games, we confirm the existence of practically relevant problems that meet the conditions in Theorem 9.
[…]
What is the potential function of a game in the class identified above? Does a Nash equilibrium exist for such a game? If yes, how could we compute it? These are the questions we examine in the next section. While answers are provided to the first two questions, we identify the challenges in answering the third one.
4 Properties of decoupled LQ potential games
We first derive the potential function for the game satisfying the conditions of Theorem 9. This potential function corresponds to a single-agent optimal control problem. However, this problem is not a standard LQ control due to the structure of the control policy arising from the decoupled state linear feedback information structure. Given that any optimizer of the potential function will be a Nash equilibrium, we verify that the control problem has an optimal solution. However, we show that characterizing or computing the best response is non-trivial.
4.1 Existence of a Nash equilibrium
Proposition 11.
The LQ game in (2) with decoupled dynamics in (5) and the decoupled state feedback information structure in (iii) under conditions (C1) and (C2) in Theorem 9 has the potential function $\Pi:\mathbb{R}^{t_{f}n}\rightarrow\mathbb{R}$:
$\Pi(\gamma)=\mathop{\mathbb{E}}_{x_{0}\sim\mathcal{D}}\big[\sum_{t=0}^{t_{f}}\bar{x}_{t}^{T}Q_{t}\bar{x}_{t}+\sum_{t=0}^{t_{f}-1}u_{t}^{T}R_{t}u_{t}\big],$ (18)
where $\bar{x}_{t}=x_{t}-d_{t}$, and the joint state, control, and decision variable are respectively, $x_{t}$, $u_{t}$, and $\gamma=k$. The loss function matrices $R_{t}\in\mathbb{R}^{N\times N}$, and $Q_{t}\in\mathbb{R}^{n\times n}$ are
$R_{t}=\begin{bmatrix}R_{t}^{11}&\dots&R_{t}^{1N}\\ \vdots&\ddots&\vdots\\ R_{t}^{N1}&\dots&R_{t}^{NN}\end{bmatrix},\ Q_{t}=\begin{bmatrix}Q_{t}^{11}&\dots&Q_{t}^{1N}\\ \vdots&\ddots&\vdots\\ Q_{t}^{N1}&\dots&Q_{t}^{NN}\end{bmatrix}$,
with $Q^{ij}_{t}$, $R_{t}^{ij}$ defined in (C1) and (C2) in Theorem 9. Note that $R_{t}$ is positive definite, and $Q_{t}$ is positive semi-definite.
{pf}
The LQ game with decoupled dynamics and the decoupled state feedback information structure in (iii) is a potential game under conditions (C1) and (C2) in Theorem 9. For this LQ game, we illustrate that the loss function of agent $i$ is a summation of equation (18) and a dummy game which is defined in [12] as a game that is independent of agent $i$ decision variables $k^{i}$, and according to Proposition 12.5 in [12], the function in equation (18) is the potential function of this LQ game.
[…]
Proposition 12.
Any LQ potential game with decoupled dynamics and the decoupled state feedback information structure in (iii) has at least one Nash equilibrium. This Nash equilibrium is the optimal structured control of the single-agent loss function identified in (18).
{pf}
The loss function is continuous and bounded below, growing to infinity whereas any $K_{t}$ grows to infinity. Thus, at least an optimal control policy exists. This policy is a Nash equilibrium as any potential function optimizer is a Nash equilibrium of the potential game. ∎
Proposition 12 does not provide insight into the computation of a Nash equilibrium. To address this, we first show that the agents’ best responses are coupled if and only if their initial states are correlated (see Proposition 13). In other words, with uncorrelated initial states, the game becomes degenerate, allowing agents to compute the optimal policies using only their local loss function and dynamics. Thus, the only non-trivial LQ potential games with decoupled dynamics and losses are the ones with correlated initial states. However, we also highlight the challenge in computing a Nash equilibrium in this case by demonstrating that dynamic programming cannot be used to compute the best response, and therefore, a Nash equilibrium of the game.
Proposition 13.
The best response of an agent $i$ in the LQ game defined in (2) with decoupled dynamics in (5) and the decoupled state feedback information structure in (iii) under conditions (C1) and (C2) in Theorem 9 depends on agent $j$’s state parameters and initial state if and only if their initial states are correlated.
{pf}
The loss function is bounded below; thus, at least one best response policy exists. If there is a unique value for which the derivative equals zero, the best response can be determined by finding the stationary point. By taking the derivative of the loss function of agent $i$ with respect to $k^{i}_{\tau}$, we obtain:
$\frac{\partial J^{i}(k^{i},k^{-i})}{\partial k^{i}_{\tau}}=2\sum_{t=\tau+1}^{t_{f}}\sum_{j=1}^{N}(b^{i})^{\top}\bigl((Q_{t}^{i})_{ij}+(k^{i}_{t})^{\top}(R^{i}_{t})_{ij}k^{j}_{t}\bigr)$
$\mathop{\mathbb{E}}_{x_{0}\sim\mathcal{D}}\left[x^{j}_{t}(x^{i}_{t,-\tau})^{\top}\right]+2\sum_{j=1}^{N}(R^{i}_{\tau})_{ij}k^{j}_{\tau}\mathop{\mathbb{E}}_{x_{0}\sim\mathcal{D}}\left[x^{j}_{\tau}(x^{i}_{\tau})^{\top}\right],$
where $x^{i}_{t,-\tau}=\prod_{t^{\prime}=0,t^{\prime}\neq\tau}^{t}(A^{i}-b^{i}k^{i}_{t^{\prime}})x_{0}^{i}$.
First, we show that the correlation of the initial states of two agents is a necessary condition for their policies to depend on each other’s parameters and initial states. If the initial state of agent $i$ is not correlated with any other agent’s initial state, it indicates that $\mathop{\mathbb{E}}\left[x_{0}^{j}(x_{0}^{i})^{\top}\right]=0$ for $j\neq i$, and $\mathop{\mathbb{E}}\left[x_{\tau}^{j}(x_{\tau}^{i})^{\top}\right]=0$ for $j\neq i$ and any $\tau\in\{0,\dots,t_{f}\}$, because $\mathop{\mathbb{E}}\left[x_{\tau}^{j}(x_{\tau}^{i})^{\top}\right]=\prod_{t^{\prime}=1}^{\tau-1}(A^{j}-b^{j}k^{j}_{t^{\prime}})\mathop{\mathbb{E}}\left[x_{0}^{j}(x_{0}^{i})^{\top}\right]\left(\prod_{t^{\prime}=1}^{\tau-1}(A^{i}-b^{i}k^{i}_{t^{\prime}})\right)^{\top}$. Hence, the best response $k^{i,*}_{t_{f}-1}$ is:
$k^{i,*}_{t_{f}-1}=\left((R^{i}_{t_{f}-1})_{ii}+(b^{i})^{\top}(Q_{t_{f}}^{i})_{ii}b^{i}\right)^{-1}(b^{i})^{\top}(Q_{t_{f}}^{i})_{ii}A^{i}.$
The other $k^{i,*}_{\tau},$ for $\tau\in\{0,\dots,t_{f}-2\}$, can be computed through dynamic programming and they do not depend on other agents’ parameters as well. Thus, the best response for agent $i$ depends only on its own parameters and initial state.
Now, we show that the correlation of the initial states of two agents is a sufficient condition for their policies to depend on each other’s parameters and initial states. For the sake of simplicity, we take $t_{f}=1$, but the same results will hold for any value of $t_{f}$. Assuming that $\mathop{\mathbb{E}}\left[x_{0}^{i}(x_{0}^{i})^{\top}\right]$ is invertible, there exists a unique stationary point and a unique $k^{i,*}_{0}$:
$k^{i,*}_{0}$= $\left((R^{i}_{0})_{ii}+(b^{i})^{\top}(Q_{1}^{i})_{ii}b^{i}\right)^{-1}\biggl\{(b^{i})^{\top}(Q_{1}^{i})_{ii}A^{i}+$
$\sum_{j\neq i}\bigl((b^{i})^{\top}(Q_{1}^{i})_{ij}(A^{j}-b^{j}k^{j}_{0})$-
$(R^{i}_{0})_{ij}k^{j}_{0}\bigr)\mathop{\mathbb{E}}\left[x_{0}^{j}(x_{0}^{i})^{\top}\right]\biggr(\mathop{\mathbb{E}}\left[x_{0}^{i}(x_{0}^{i})^{\top}\right]\biggl)^{-1}\biggr\}.$
If the correlation between the initial states of agent $i$ and at least one other agent $j$ is non-zero, the optimal $k^{i,*}_{0}$ depends on the value of $A^{j}$, $b^{j}$, $k^{j}$, and $x^{j}_{0}$. This result holds even for non-potential, decoupled LQ games. ∎
The proof above can be generalized to show that the computation of the optimal decoupled linear control policy of an agent $i$ using dynamic programming is not possible if its initial state is correlated with the initial state of at least one other agent. In particular, the best response $k^{i,*}_{t_{f}-1}$ depends on the values of $x_{t_{f}-1}^{i}$, which in turn depends on the policies at the previous time steps. This interdependence precludes the use of the dynamic programming method to compute the best response.
4.2 Challenge in computing the Nash equilibria
Recent works have proven that policy gradient can be applied for computing the optimal control policy in LQ problems [8] and LQ problems with sparsity constraints [9]. For the LQ regulator problem, [8] proved that policy gradient converges to the unique optimum, based on the uniqueness of the global optimum and the gradient-dominated and almost smooth nature of the loss function (see Lemmas 3 and 6 in [8]). In [9], this result was extended to finite-horizon structured control problems, a class that encompasses our decoupled LQ potential games. However, the convergence proof in [9] relied on the so-called quadratic invariance property of the sparsity constraints. Unfortunately, as shown below, the quadratic invariance is not generally satisfied for the decoupled information structure considered here.
Counterexample for the quadratic invariance: A subspace $\mathcal{K}$ is quadratic invariant with respect to a matrix $\textbf{CP}_{12}$ if and only if
$\textbf{KCP}_{12}\textbf{K}\in\mathcal{K}\text{, }\qquad\forall\textbf{K}\in\mathcal{K},$
where C and P are matrices depending on the system parameters.
We provide an example of a decoupled LQ game with two agents and time horizon $t_{f}=2$ that does not respect the condition of quadratic invariance as presented in [9, Definition 1]. The notation used here is aligned with that of [9]. For simplicity, we choose the system matrices $A$ and $B$ as the identity matrix $A=B=I_{2}.$ We also assumed perfect knowledge of the state, thus the matrix $C$ is the identity matrix as well.
The subspace $\mathcal{K}$ must satisfy the sparsity constraint due to the information structure outlined in (iii). In our scenario, $\mathcal{K}=\text{diag}(k_{1},k_{2},k_{3},k_{4})$ with $k_{1},k_{2},k_{3},k_{4}\in\mathbb{R}$. Given the matrix $\textbf{K}=I_{4}\in\mathcal{K}$, we have that $\textbf{KCP}_{12}\textbf{K}=\begin{bmatrix}&0_{2\times 2}&0_{2\times 2}\\ &I_{2}&0_{2\times 2}\end{bmatrix}\notin\mathcal{K}.$ Thus, the subspace $\mathcal{K}$ is not quadratic invariant with respect to matrix $\textbf{CP}_{12}$. Note that the issue here is that decoupled policies are memoryless. Namely, if we allow the policy of agent $i$ at time $t$ to depend on its state at previous times the same information structure can result in quadratic invariance.
[…]
4.3 Convergence of policy gradient to a stationary point
Here, we derive conditions for convergence to a stationary point of the policy gradient. Consider the joint decision-variables $k=(k^{1},k^{2},...,k^{N})\in\mathbb{R}^{t_{f}nN}$, where $k^{i}=(k^{i}_{1},...,k^{i}_{t_{f}-1})$. Starting from an arbitrary $k^{i}(0)\in\mathbb{R}^{t_{f}n}$, each agent updates its feedback policy as
$k^{i}(m+1)=k^{i}(m)-\eta(m)\lambda^{i}\nabla_{k^{i}}J^{i}(k(m)).$ (19)
The step size of each agent is composed of two terms. The first one, $\eta(m)\in\mathbb{R}_{>0}$, is the same for all agents, while the second term, $\lambda^{i}\in\mathbb{R}_{>0}$, can be chosen independently by each agent. From the definition of derivative and equation (7), it follows that the gradient of the loss function $J^{i}(k)$ with respect to $k^{i}$ is equal to the gradient of the potential function $\Pi(k)$ defined in (18) with respect to $k^{i}$. The update iteration (19) is equivalent to:
$k^{i}(m+1)=k^{i}(m)-\eta(m)\lambda^{i}\nabla_{k^{i}}\Pi(k(m)).$ (20)
Stacking the iterations in (20) for all agents we obtain:
$k(m+1)$ $=k(m)-\eta(m)\Lambda\nabla\Pi(k(m)),$ (21)
$\Lambda$ $=\text{blockdiag}(\lambda_{1}I_{t_{f}n^{1}},...,\lambda_{N}I_{t_{f}n^{N}}).$
Note that the term $\Lambda\nabla\Pi(.)\in\mathbb{R}^{t_{f}n}$ is a descent direction, $(\nabla\Pi(.))^{T}\Lambda\nabla\Pi(.)<0,$ and $\Pi$ is continuously differentiable and bounded below in every direction. Thus, from Lemma 3.1 in [19], there exists an $\eta$ satisfying the Wolfe conditions, and it can be found using the bisection algorithm presented in [3, Proposition 5.5]. It follows from [19] that $\sum_{m\geq 0}\biggl(\frac{\nabla\Pi(k(m))^{T}\Lambda\nabla\Pi(k(m))}{\|\Lambda\nabla\Pi(k(m))\|}\biggr)^{2}<\infty.$ This condition is true if and only if $\nabla\Pi(k_{m})$ converges to 0, which implies that: $\lim_{m\to\infty}k(m)=k^{s},$ where $k^{s}$ is a stationary point. Note that variants of the policy gradient above, such as stochastic policy gradient assuming no access to the model, can be developed.
4.4 Simulation
In this subsection, we implement the policy gradient method introduced in equation (21). Our goal is to evaluate the convergence of the policy gradient to a stationary point.
We consider a game with $N=4$ agents, where each agent has a state of dimension $n^{i}=1$ and scalar control actions. The time horizon is $t_{f}=10$. To construct the example, we randomly generate the parameters of the game. The values of $A^{i}\in\mathbb{R}$ and $b^{i}\in\mathbb{R}$ are independently sampled from a normal distribution with zero mean and unit variance. Since the matrices $Q_{t}\in\mathbb{R}^{4\times 4}$ and $R_{t}\in\mathbb{R}^{4\times 4}$ of the potential function should be respectively positive semidefinite and positive definite, we first generate the auxiliary matrices $M_{t}\in\mathbb{R}^{4\times 4}$ and $N_{t}\in\mathbb{R}^{4\times 4}$ with entries drawn from the same normal distribution, and then we set $Q_{t}=M_{t}^{\top}M_{t}$ and $R_{t}=N_{t}^{\top}N_{t}+0.01I_{4}$.
Starting from a random $k(0)$, we run the policy iteration for $10^{4}$ steps. We observe that $k(m)$ converges to a stationary point $\tilde{k}$. Then, we repeat the experiment twenty times with different random initializations $k(0)$. The values $k^{i}_{t}(0)$ are randomly sampled from a standard normal distribution. As shown in Fig 1(a), all simulations converge to a stationary point. Furthermore, Fig 1(b) shows that the distance from the previously computed $\tilde{k}$ approaches zero in all simulations, indicating convergence to the same stationary point. As proved in Proposition 12, the game admits at least a Nash equilibrium. Since we found only one stationary point, and any Nash equilibrium has to be a stationary point, it would seem plausible that the $\tilde{k}$ is a Nash equilibrium. However, as discussed in Subsection 4.2, proving that this point is a Nash equilibrium requires demonstrating that no agent can improve their outcome by unilaterally selecting a different policy. Due to the nonlinear dependence of the optimal strategy $k^{i,*}_{t}$ on the policies at both previous and successive steps, checking for potential improvements would necessitate an exhaustive search over the entire policy space. Specifically, this involves exploring a space with dimensions $n\times t_{f}=40$ in our current setting.
(a) $|\nabla\Pi(k(m))|$
(b) |k(m)k~||k(m)-$\tilde${k}|
Figure 1: Convergence of policy gradient with the update rule in equation (21) to a stationary point $\tilde{k}$ for 20 simulations. In these figures, $m$ indicates the iteration number, and the abscissa and the ordinate are on a logarithmic scale.
5 Conclusion
Motivated by the desirable properties of dynamic potential games, our work focused on advancing a fundamental understanding of linear quadratic potential games. To this end, we considered finite-horizon linear quadratic games and derived conditions under which the game would be potential. Notably, we showed that an example of a general class of scalar linear quadratic games with full-state feedback is potential if it deviates slightly from an identical interest game. Additionally, we investigated a subclass of games with decoupled dynamics and decoupled state feedback, revealing the existence of potential games beyond identical interest games. Furthermore, we analyzed the potential function of this subclass and demonstrated the existence of at least one Nash equilibrium. We illustrated how initial state correlations can influence agents’ behavior and highlighted the challenge of computing a Nash equilibrium policy.
Future research can explore approaches to compute the Nash equilibria of the identified potential LQ games, investigate alternative control structures beyond the decoupled state linear feedback, and extend the potential LQ game characterization to an infinite-horizon setting.
{ack}
Sara Hosseinirad is financially supported by the Natural Sciences and Engineering Research Council of Canada. Giulio Salizzoni is financially supported by the Swiss National Science Foundation, grant number 207984.
Appendix A Additive Noise
Here, we prove that the results in Proposition 5 and Theorem 9 can be generalized to the LQ game with noise. For this purpose, we introduce the loss function of the game with noisy dynamics, compare the derivatives in condition (10) of Lemma 4 for this loss function and the one without noise, and illustrate that the final potential conditions in Proposition 5 and Theorem 9 are the same for both loss functions. We assume the dynamics of the game include noise as follows.
$x_{t+1}=Ax_{t}+\sum_{i=1}^{N}B^{i}u^{i}_{t}+w_{t},\quad x_{0}\sim\mathcal{D},\ w_{t}\sim\mathcal{W}$ (22)
where $w_{t}$ is a random noise that is assumed to be drawn from the distribution $\mathcal{W}$ with zero mean. The loss function will be as follows.
$J^{i}(\gamma)=$ $\mathop{\mathbb{E}}_{x_{0}\sim\mathcal{D}}\bigl[\sum_{t=0}^{t_{f}-1}\mathop{\mathbb{E}}_{w_{t}\sim\mathcal{W}}c_{t}(x_{t+1},u_{t})\bigr],$ (23)
where the stage loss $c_{t}(x_{t+1},u_{t})$ is defined as follows.
$c_{t}(x_{t+1},u_{t}):=$ $(x_{t+1}-d_{t+1})^{T}Q_{t+1}^{i}(x_{t+1}-d_{t+1})$
$+\sum_{j=1}^{N}\sum_{h=1}^{N}(u_{t}^{j})^{T}(R^{i}_{t})_{jh}u^{h}_{t},$ (24)
Note that for decoupled dynamics, we add a random noise $w^{i}_{t}$ drawn from the distribution $\mathcal{W}^{i}$ with zero mean to the set of decoupled dynamics in equation (5). By defining $w_{t}:=(w^{1}_{t},\dots,w^{N}_{t})$, the loss function of the LQ game with noisy decoupled dynamics can be written as the one in equation 23. Considering the loss function defined above, the definitions of Nash equilibrium, potential game, identical interest games for LQ games, (Definitions 12, and 3) the game pseudo-gradient, $\mathcal{G}(\gamma)$, and its Jacobian in equation (9) do not change. Hence, Lemma 4 is still valid for the game with noisy dynamics. Note that the derivative of $c_{t}(x_{t+1},u_{t})$ with respect to $x_{t}$ or $u_{t}^{j}$ is similar to the game without noise. Therefore, we only investigate the derivative of $x_{t+1}$ and $u_{t}^{i}$ with respect to the decision variable $\gamma_{\tau}^{j}$ in the presence of the noise.
Proposition 5 part (a): For the open-loop information structure, note that the decision-making variable is $\gamma_{t}^{j}=u_{t}^{j}$, and the game dynamics can be written as $x_{t+1}=a^{t}x_{0}+\sum_{\tau=0}^{t}a^{t-\tau}[b^{1}u^{1}_{\tau}+b^{2}u^{2}_{\tau}+w_{\tau}]$. Thus, the derivative of $x_{t+1}$ and $u_{t}^{i}$ with respect to $u_{\tau}^{j}$ is as follows.
$\frac{\partial x_{t+1}}{\partial u_{\tau}^{j}}=$ $\begin{cases}A^{t-\tau}B^{i}&t\geq\tau\\ 0&t<\tau\end{cases}$
$\frac{\partial u_{t}^{i}}{\partial u_{\tau}^{j}}=$ $\begin{cases}1&t=\tau\ \text{and}\ i=j\\ 0&\text{otherwise}\end{cases}$
The derivatives above are similar to the ones without noise. This proves that the conditions in Proposition 5, part (a), are valid for the game with noise.
Proposition 5 part (b): For the full-state feedback information structure, note that the decision-making variable is $\gamma_{t}^{j}=K_{t}^{j}$. The derivative of $x_{t+1}$ and $u_{t}^{i}$ with respect to $K_{\tau}^{j}$ can be derived using the chain rule and assuming the states are scalar (as mentioned in Example 3) as follows.
$\frac{\partial x_{t+1}}{\partial K_{\tau}^{j}}=$ $\begin{cases}-\prod_{t^{\prime}=\tau+1}^{t}\frac{\partial x_{t^{\prime}+1}}{\partial x_{t^{\prime}}}b^{j}x_{\tau}&t>\tau\\ -b^{j}x_{t}&t=\tau\\ 0&t<\tau\end{cases}$
$\frac{\partial u_{t}^{i}}{\partial K_{\tau}^{j}}=$ $\begin{cases}-K_{t}^{i}\frac{\partial x_{t}}{\partial K_{\tau}^{j}}&t>\tau\\ -x_{t}&t=\tau\ \text{and}\ i=j\\ 0&\text{otherwise}\end{cases}$
Note that $\frac{\partial x_{t+1}}{\partial x_{t}}$ is not a function of $w_{t}$; therefore, the derivatives above are similar to the ones without noise. This proves that the conditions in Proposition 5, part (b), are valid for the game with noise.
Theorem 9: For the decoupled state feedback information structure, note that the decision-making variable is $\gamma_{t}^{j}=k_{t}^{j}$. The derivative of $x^{i}_{t+1}$ and $u_{t}^{i}$ with respect to $k_{\tau}^{i}$ can be derived using the chain rule as follows.
$\frac{\partial x^{i}_{t+1}}{\partial k_{\tau}^{i}}=$ $\begin{cases}-\prod_{t^{\prime}=\tau+1}^{t}\frac{\partial x^{i}_{t^{\prime}+1}}{\partial x^{i}_{t^{\prime}}}b^{i}(x^{i}_{\tau})^{T}&t>\tau\\ -b^{i}(x^{i}_{t})^{T}&t=\tau\\ 0&t<\tau\end{cases}$
$\frac{\partial u_{t}^{i}}{\partial k_{\tau}^{i}}=$ $\begin{cases}-k_{t}^{i}\frac{\partial x^{i}_{t}}{\partial k_{\tau}^{i}}&t>\tau\\ -(x^{i}_{t})^{T}&t=\tau\\ 0&t<\tau\end{cases}$
Note that the derivative of $x^{i}_{t+1}$ and $u_{t}^{i}$ with respect to $k_{t}^{j}$ for $j\neq i$ are zero, and $\frac{\partial x^{i}_{t+1}}{\partial x^{i}_{t}}$ is not a function of $w_{t}$. Hence, the derivatives above are similar to the ones without noise. These results indicate that the conditions derived for the game with the decoupled dynamic defined in equation (5) and a decoupled state feedback information structure in (iii) to be potential do not depend on noise, and consequently, adding the noise does not affect the results.